OpenStudy (sf_r):

Pre-Algebra Help! Failing currently, i wanna do this thing without doing something wrong.

OpenStudy (sf_r):

Hello!

OpenStudy (sf_r):

SO, i have 52% Currently.

OpenStudy (sf_r):

And... I need to bring 70% so i have to do it right for that.

OpenStudy (sf_r):

I will just put out the question anyway: What is the volume of a cube with a side length of 15 inches? A. 45 in3 B. 3375 in3 C. 225 in3 D. 1350 in3

OpenStudy (usukidoll):

first you need the volume of a cube formula

OpenStudy (sf_r):

True, i don't have one actually! Let me check.

OpenStudy (sf_r):

Well i don't have one ;-;

OpenStudy (sf_r):

I hate my online schoo0o0o0o0ol.

OpenStudy (usukidoll):

V=a^3 where a = is the side length . since a = 15 V = 15 x 15 x 15 go compute that :)

OpenStudy (sf_r):

Okay let me take a look you are a genius anyway :D

OpenStudy (sf_r):

So it's 3375 in

OpenStudy (usukidoll):

ya

OpenStudy (sf_r):

Wow you are amazing..!!!

OpenStudy (sf_r):

Well i know about the variable but i don't focus much on it!

OpenStudy (sf_r):

I just need this assessment done hopefully i'll make some i mean you cause of the help will make some progress for me!

OpenStudy (usukidoll):

feel free to donate some owlbucks to me ;)

OpenStudy (sf_r):

Okay i will see what is that cause i am a noob ;-; but i will if ik how hehe.

OpenStudy (usukidoll):

it's like a tip jar ;)

OpenStudy (sf_r):

Is that answer for this 1942? How many milliliters equal 1942 cm3? A. 1.942 B. 19,420 C. 1942 D. 194.2

OpenStudy (sf_r):

Yeah i will :D

OpenStudy (usukidoll):

looks like conversion like from mililiters to centimeters cube. hold on a sec

OpenStudy (sf_r):

yes it is my middle unit assessment so this will include 2 or 3 topics

OpenStudy (usukidoll):

it is 1942

OpenStudy (sf_r):

Alright and cn you tell me the way to tip you that owl thing and medal please!

OpenStudy (usukidoll):

check your pm

OpenStudy (sf_r):

What is the volume of the cylinder? A. 60π ft3 B. 720π ft3 C. 300π ft3 D. 600π ft3

OpenStudy (sf_r):

Okay dokie!

OpenStudy (usukidoll):

alrighty so we need that formula pi × radius^2 × height height = 12 radius =5

OpenStudy (usukidoll):

pi x 5^2 x 12 pi x 25 x 12=300pi

OpenStudy (sf_r):

Alright so we will solve this for volume?

OpenStudy (usukidoll):

mhm

OpenStudy (sf_r):

Okay but i don't know about the PI we're you double like 25235235B times. ;-;

OpenStudy (usukidoll):

pi is the same as π it was already calculated and I got 300 pi or 300 π

OpenStudy (sf_r):

oh i am dumb

OpenStudy (sf_r):

I think so they are 3 more, it's like 11:09 here and i have to do this thinggg

OpenStudy (usukidoll):

uh oh T_T

OpenStudy (sf_r):

The base area of a prism is 140 square feet. Its height is 6 feet. What is the volume of the prism? A. 840 ft3 B. 280 ft3 C. 146 ft3 D. 420 ft3 Well for his it'll be easy if you have an image i think so

OpenStudy (sf_r):

I will try this 1 as well.

OpenStudy (usukidoll):

don't you need the volume of the prism formula????

Directrix (directrix):

> How many milliliters equal 1942 cm3? On the second question, it appears that you are converting linear measure to volume measure. Is the question written correctly? Just wondering.

OpenStudy (sf_r):

Should i try that question again if you want me too!?!?

Directrix (directrix):

What is the volume of the prism? V = area of base times height V = 140*6 =

OpenStudy (usukidoll):

^ that's what I got XD

OpenStudy (sf_r):

Woah Y'all are like my teachers :D

OpenStudy (sf_r):

And these letters are called variable in math right?

OpenStudy (usukidoll):

yeah like x y z

Directrix (directrix):

>1942 I don't see how this could be the answer to that second question. I think that the question is posted without the complete text.

OpenStudy (usukidoll):

odd a conversion calculator will have it at 1942 ? ~_~

OpenStudy (sf_r):

Let me see.

OpenStudy (sf_r):

For the second one this is the whole text: How many milliliters equal 1942 cm3? A. 1.942 B. 19,420 C. 1942 D. 194.2

OpenStudy (sf_r):

What is the approximate volume of the cone? Use 3.14 for π. A. 1272 cm3 B. 4239 cm3 C. 2120 cm3 D. 6359 cm3 Um @Usuki i forgot to solve the first problem actually ;-;

OpenStudy (sf_r):

For this i randomly picked " 6359 cm3" which is totally wrong i have a feeling.

OpenStudy (usukidoll):

V = 1/3 b x h b = base h = height b = 15 h = 9 whoa that's a bit small maybe V = 1/3r^2h pi h = 9 r = 15

OpenStudy (sf_r):

lol You're so amazing! Wish me luck i pass i'll come to tell you if i :D

OpenStudy (sf_r):

Oh my god you can go if you want for your work.

OpenStudy (sf_r):

Exactly i don't know how to do this 1 specially: The base area of a prism is 140 square feet. Its height is 6 feet. What is the volume of the prism? A. 840 ft3 B. 280 ft3 C. 146 ft3 D. 420 ft3

OpenStudy (sf_r):

I posted this before but you said you'll need something that will help or something basic idk

OpenStudy (sf_r):

This is 840 I think

OpenStudy (sf_r):

I will go for 840 WISH this is right!!!

OpenStudy (sf_r):

What is the capacity of the cube in liters? Mid-Unit Test A. 3.43 L B. 343 L C. 0.343 L D. 343,000 L

OpenStudy (sf_r):

I GOT A MASTERY :D 80+ :DDDD

OpenStudy (sf_r):

I have to go....

OpenStudy (sf_r):

Aye brian! If usuki is back say to her/him ;-; thanks for the help (if you can)

OpenStudy (usukidoll):

woo congrats @SF_R !!!

OpenStudy (sf_r):

This is still on??

OpenStudy (kdr1308):

@SF_R I have the same problem with online school, you can go to Khan Academy, it helps

OpenStudy (sf_r):

Yeah i watch there videos they are helpful! But they are some problems that i solve and i get no where...

OpenStudy (kdr1308):

Ah, I see.. well we all are here to help!

OpenStudy (sf_r):

Yes, right now i am doing science well i am good at it hehe, 7th grade and.. but

OpenStudy (sf_r):

I am failing math

OpenStudy (sf_r):

omg im dumb.

OpenStudy (sf_r):

anyways i am a noob here and i see this coin and it's increasing i have 3 of them.

alones (alones):

Don't cal your self dumb LOL

OpenStudy (sf_r):

dunno actually what i am ;-;

OpenStudy (sf_r):

Can someone give me a good example of "What are the 3 challenges of life" :P